# GraphMend: Code Transformations for Fixing Graph Breaks in PyTorch 2

### v1

@@ IV-C Jac Compilation Pipeline
foreach Dynamo Entry Point $r$ in UniiR do
  Initialize worklist with successors of $r$
  while worklist not empty do
     Pop node $n$ from worklist; push unseen successors of $n$
@@
     if $n$ is an IfStmt then
        $E\leftarrow$ subexpressions of $\texttt{cond}(n)$
        $dyn\leftarrow\textbf{false}$
        foreach $e\in E$ do
           if $e$ has a torch attribute then
@@
              if TorchAttrTable[$a$] = dynamic then
                 $dyn\leftarrow\textbf{true}$; break
              end if
           end if
        end foreach
       if not dyndyn then
           if $E$ is input/tensor-value dependent (via SymbolTable + CFG) then
              $dyn\leftarrow\textbf{true}$
           end if
        end if
       if dyndyn then
           Tag $n$ as GraphBreak[DynCtrlFl]
        end if
     end if
@@
     if $n$ is a call to print or logger.* then
        Tag $n$ as GraphBreak[logger/print]
     end if
  end while
end foreach
@@ IV-D GraphMend Implementation
IV-D1 Dynamo Entry Point Analysis Pass
To identify potential TorchDynamo compilation entry points, this pass traverses the AST and analyzes each node. A node is tagged when a function or an nn.Module object wrapped or decorated with torch.compile is encountered.
If no occurrences of torch.compile are found in the codebase, GraphMend compilation is skipped. Otherwise, the tagged nodes serve as entry points for subsequent passes.
IV-D2 Graph Break Type Analysis Pass
We implement a compound analysis pass on UniiR to identify potential graph-break locations, following the procedure in Algorithm 1. Starting from each Dynamo entry point, the pass traverses the control-flow graph (CFG) to locate program points that may induce graph breaks. For conditional branches, the pass inspects the subexpressions of the condition. If any involve Torch attributes known to be dynamic (e.g., .sum()), the branch is marked as dynamic control flow. Otherwise, a use-def analysis over the symbol table and CFG determines whether the condition depends on model inputs or tensor values; if so, it is also tagged as dynamic. In addition, side-effecting operations such as print and logger.* calls are detected and annotated as logging-induced graph breaks as in Algorithm 1.
These annotations guide the subsequent transformation passes that eliminate graph breaks.
IV-D3 AST Transformation Pass
After identifying graph-break locations, the AST must be modified to eliminate them. As discussed earlier in this section, these fixes can be applied through source-level transformations. Following the same approach, the AST is modified according to the graph-break tags identified in the previous pass. Based on these tags, one of the following transformations is performed.
@@ IV-D3 AST Transformation Pass
To address data-dependent control-flow graph breaks, we first evaluate the branch condition and store it as a predicate. The control flow is then rewritten using torch.where, which enables predicated execution of both branches. This process is illustrated in Figure 6. In the AST Transformation Pass (Figure 6), the original branch if x.sum() > 5: ... else: ... is reduced by extracting the condition into a symbolic variable pred (R 1). Next, the two branch assignments are merged into a single dataflow expression using torch.where(pred, e1, e2). Finally, the if and else AST nodes are removed, and their child nodes are reconnected to the parent, producing a transformed tree in which both execution paths are explicitly represented.
Graph-Epilogue Deferred Side Effect Transformation
For print or logger calls, the message is first stored in a temporary variable. This variable is then passed to the outermost forward function and executed at the end of the function (epilogue) before returning, as shown in Figure 6. In the example ASTs in Figure 6, the original call print("Logger message") is rewritten as a deferred assignment deferred = "Logger message". This assignment is then moved to the epilogue block, where it is printed after graph execution (R 2). In this way, the logging effect is preserved while remaining safely outside the traced region. Note that any computation in the return statement must be moved before the epilogue, stored in a temporary variable, and then returned afterward to avoid introducing a graph break.
IV-E Healing UniiR and Downstream Flow
After the AST transformations, the UniiR is no longer consistent. Therefore, the symbol table and CFG must be rebuilt using the same passes that originally created them. As shown in Figure 6, once the GraphMend pass is complete, the SymTabBuildPass and CFGBuildPass are rerun to restore UniiR for downstream compilation.
@@ IV-E Healing UniiR and Downstream Flow
TABLE I: Models in benchmark suite
biogpt [15] 2 logger calls 408
blenderbot-400M-distill [16] 3 logger calls 489
flan-t5-large [17] 3 logger calls 2130
longformer-base-409 [18] 5 logger calls, tensor.item() call 913
moe-minicpm-x4-base [19] 15 Dynamic shape operator 280
Phi-4-mini-instruct [20] 5 Dynamic control flow 1541
Qwen-Audio-Chat [21] 2 Dynamic control flow 1647
tiny-random-PegasusForCausalLM [22] 2 logger calls 64
V Evaluation
Experimental Setup
We implement GraphMend on top of Jaseci [6, 7, 9], an open-source research infrastructure that provides both compiler and runtime control as a Python superset. Our experiments are conducted on NVIDIA RTX 3090 and A40 GPUs, using TorchInductor as the backend compiler.
Benchmark Suite
To construct our benchmark, we (1) randomly sample 65 Hugging Face [10] models spanning diverse architectures and task categories; (2) identify models that trigger FX graph breaks; and (3) retain the eight models with breaks for evaluation. Table I summarizes the resulting suite, including the number and causes of graph breaks and the total operator count. A similar methodology was used in prior works to construct benchmark suite [2].
Profiling Methodology
We profile the forward pass using PyTorch Profiler [23]. For each model, we evaluate two configurations: original model and graph breaks fixed version. Each configuration is executed for 7 iterations: one cold-start iteration followed by six warm iterations. The profiler records CPU and CUDA activities and operator-level events; traces are exported and inspected with Chrome’s trace viewer.
V-A GraphMend’s Ability to Fix Graph Breaks
TABLE II: Graph break counts in the original model and fix rates achieved by applying GraphMend across the benchmark suite.
@@ V-A GraphMend’s Ability to Fix Graph Breaks
First, we check GraphMend’s ability to fix graph breaks across the benchmark suite. Table II shows that GraphMend fully eliminates graph breaks in 6 of the 8 models. It partially fixes longformer-base-4096 and leaves moe-minicpm-x4-base unchanged. By cause, GraphMend removes all breaks due to logging statements and data-dependent control flow. The remaining breaks: tensor.item() and dynamic-shape operators could not be fixed by code transformations. For the performance analysis that follows, we report results on the seven models whose break counts were fully eliminated or reduced, excluding moe-minicpm-x4-base.
V-B Latency Improvement per Forward Pass Using GraphMend to Fix Graph Breaks
We measure Relative Latency Improvement (%) as
$100\times\frac{T_{\text{break}}-T_{\text{fixed}}}{T_{\text{break}}},$
where $T_{\text{break}}$ is the per-forward-pass latency for the benchmarked model with graph breaks, and $T_{\text{fixed}}$ is the latency after applying GraphMend fixes.
We report the relative improvement per forward pass for break-fixable models in the benchmark suite, using PyTorch Profiler data for both cold and warm (steady-state) runs.
Cold Start Latency Improvement
Figure 7(a) reports results on both RTX 3090 and A40 GPUs. Across models, we observe a relative latency reduction of about 30–75% for both GPUs, indicating substantial gains in cold runs. At cold start, each graph break incurs significant overhead because every CUDA graph requires a separate record and caching. After eliminating these breaks, only a single (or fewer) CUDA Graph capture is needed, resulting in fewer records and reduced capture time, which minimizes overhead. Consequently, we achieve larger relative latency improvements per forward pass in cold runs. A more detailed analysis is provided in  § V-D.
Steady-State Latency Improvement
Figure 7(b) shows steady-state improvements on RTX 3090 and A40 GPUs, ranging from about 2.5% to 25% across models. After the cold run, the captured graph is cached and replayed, removing the high capture overhead present initially. However, when graph breaks persist, execution still incurs GPU–CPU switches (as discussed in  § V-D), which add latency. Eliminating these breaks with GraphMend removes such switches, yielding additional improvements in relative latency per forward pass.
Notably, tiny-random-PegasusForCausalLM shows over 25% improvement. Since this model is very small, with only 64 operators, the fixed costs from graph breaks (such as host–device synchronization, and extra graph launches) take up a large portion of the runtime. Removing these breaks therefore gives a huge speedup per forward pass.
@@ V-B Latency Improvement per Forward Pass Using GraphMend to Fix Graph Breaks
V-C Relative Throughput Improvement by Fixing Graph Breaks
We measure Relative Throughput Improvement (%) as
$100\times\frac{Tp_{\text{fixed}}-Tp_{\text{break}}}{Tp_{\text{break}}},$
where $Tp_{\text{break}}$ is the throughput (output tokens per second) for the model with graph breaks, and $Tp_{\text{fixed}}$ is the throughput after applying GraphMend fixes.
We measure the relative throughput across the models in the benchmark suite that are fixable through GraphMend for both RTX 3090 and A40. Figure 7(b) shows a relative throughput improvement of about 5-8% across all models on both GPUs. This improvement follows from the relative latency reduction per forward pass observed in Section V-B: as each step becomes faster, the end-to-end tokens per second increase. The variation of improvement across models highlights an important point: the magnitude of improvement depends on the extent to which a graph break affects the forward path. When the impact of a break is small, the corresponding throughput gain is also limited.
For example, in both Phi-4-mini-instruct and Qwen-Audio-Chat, the root cause of the breaks is dynamic control flow. We observe a 7.5%–8% improvement for Qwen-Audio-Chat, compared to 5%–6% for Phi-4-mini-instruct from both GPUs. The difference arises because the break in Phi-4-mini-instruct occurs near the beginning of the forward function, whereas in Qwen-Audio-Chat it occurs in the middle, where its impact is higher. Thus, the location of a break determines its effect: fixing some breaks yields larger throughput gains, while fixing others has a smaller impact.
V-D Overhead Analysis of Graph Breaks
We conducted an analysis to evaluate how fixing graph breaks with GraphMend reduces overhead. We analyze Qwen-Audio-Chat model running on an A40 GPU which has 2 graph breaks.
@@ V-D Overhead Analysis of Graph Breaks
Figure 10: Profiler tracer of cold run of Qwen-Audio-Chat model run on A40 GPU.
@@
To ensure a consistent comparison in this section, all activity plots (Figures 11(a), 11(b), 11(c), and 11(d)) are drawn only for the time interval between the first and second CUDA graph executions in the original model; for the fixed model, we consider the identical time region.
@@
Figure 9(a) shows the profiler trace for the original model. Execution is split into three CUDA graphs because of two graph breaks, with idle GPU periods between the first and second FX graph executions. This is clearly visible in Figure 11(a), the GPU becomes idle while the CPU handles control flow in eager mode, including a device-to-host (D2H) memcpy triggered by dynamic control flow (graph break reason for this model). This CPU fallback introduces an overhead.
After applying GraphMend, the rewritten code replaces the dynamic branch with a GPU-supported torch.where. The resulting trace (Figure 9(b)) shows one continuous CUDA graph. As confirmed in Figure 11(c), GPU activity remains uninterrupted and CPU fallbacks disappear. Thus, GraphMend eliminates the overhead of CPU–GPU context switching during steady-state runs.
@@
Cold runs introduce even larger overheads. In the profiler trace of original model (Figure 10), each graph break triggers a separate CUDA Graph recording and caching, leading to long idle GPU phases and heavy CPU involvement as shown in Figure 11(b). In this interval, we observe that GPU idle time dominates, while the CPU is heavily active, reflecting significant overhead. By contrast, the fixed model (Figure 11(d)) executes with a single capture, avoiding repeated records and reducing idle GPU time. Overall, GraphMend reduces overhead more significantly in cold runs than in steady state, explaining the substantial cold-start latency reductions per forward pass reported in Figure 7(a).
V-E GPU Runtime Kernel Statistics Analysis
We analyze how fixing graph breaks affects GPU kernel counts, ordering, and scheduling, and whether this enables new optimization opportunities. Using the Phi-4-mini-instruct model on an A40 GPU, the profiler reports 404 executed kernels for the original model and 393 for the fixed model with a reduction of 11. This drop indicates more fusion and fewer launches, as removing breaks allows the compiler to capture larger regions and emit fewer, larger kernels.
Figure 12 illustrates this change. In the original model, 3 small kernels appear that are replaced in the fixed model by a single fused kernel. We also observe reordering: one kernel from the original run shifts upward in the fixed run. With one CUDA Graph, the runtime can reorder independent regions and pack launches more effectively, reducing stalls and improving locality.
Graph breaks force separate CUDA Graph captures, blocking fusion across the break and restricting scheduling to smaller windows. Removing them allows larger portions of the forward pass to be captured together, enabling the compiler and runtime to fuse operations, remove redundant launches, and reorder independent work. As a result, kernel gaps shrink and device utilization improves.
@@ V-E GPU Runtime Kernel Statistics Analysis
VI Related Work
PyTorch has long sought to balance performance with Python’s flexibility. TorchScript’s tracing and scripting enabled ahead-of-time graph capture but struggled with dynamic control flow and unsupported constructs [12, 11]. PyTorch 2 introduced TorchDynamo for bytecode-level capture and TorchInductor for backend compilation [2, 1], yet graph breaks remain common when encountering data-dependent branches or Python I/O. Our work reduces these breaks by rewriting source code before execution, allowing more continuous FX graphs.
General compiler infrastructures such as XLA [24], TVM [25], and MLIR [26] provide IR-based optimization, while Glow [27], TASO [28], and ONNX Runtime [29] extend portability across hardware and frameworks. These systems assume large graphs are already available, whereas we restructure PyTorch code to maximize graph continuity so their optimizations can be applied more effectively.
Dynamic models motivate systems like GRAPE [30] and Triton [31, 32], which manage runtime variability through efficient kernel generation. PyTorch’s compiler stack still falls back to eager execution in such cases, while our method removes common break sources by rewriting branches into tensorized forms that remain graph-compatible.
Program transformations more broadly, as seen in Halide [33], Lantern [34], and DLVM [30], show the potential of semantics-preserving rewrites for optimization. Our approach follows this tradition but targets PyTorch specifically, introducing lightweight transformations that enable TorchDynamo to capture larger uninterrupted graphs.
VII Conclusion
We presented GraphMend, a source-level compiler technique for PyTorch that systematically eliminates common FX graph breaks before execution. By restructuring code constructs such as dynamic control flow and Python I/O operations, GraphMend reduces fragmentation in the compilation pipeline and enables larger unified graphs to be captured by TorchDynamo. Our evaluation across a diverse suite of Hugging Face models demonstrates that GraphMend removes nearly all fixable breaks and consistently improves runtime efficiency, with substantial latency and throughput gains on modern GPUs.

### v2

@@ 4.1 Jac Compilation Pipeline
1
2foreach Dynamo Entry Point $r$ in UniiR do
3   Initialize worklist with successors of $r$
4 while worklist not empty do
5      Pop node $n$ from worklist; push unseen successors of $n$
6
@@ 5      Pop node $n$ from worklist; push unseen successors of $n$
7    if $n$ is an IfStmt then
8       $E\leftarrow$ subexpressions of $\texttt{cond}(n)$
9       $dyn\leftarrow\textbf{false}$
10       foreach $e\in E$ do
11          if $e$ has a torch attribute then
@@ 9       $dyn\leftarrow\textbf{false}$
12             if TorchAttrTable[$a$] = dynamic then
13                $dyn\leftarrow\textbf{true}$; break
14
15
16
17       if not dyndyn then
18          if $E$ is input/tensor-value dependent (via SymbolTable + CFG) then
19             $dyn\leftarrow\textbf{true}$
20
21
22       if dyndyn then
23            Tag $n$ as GraphBreak[DynCtrlFl]
24
25
26
@@ 23            Tag $n$ as GraphBreak[DynCtrlFl]
27    if $n$ is a call to print or logger.* then
28         Tag $n$ as GraphBreak[logger/print]
29
30
31
@@ 4.2 GraphMend Implementation
Dynamo Entry Point Analysis Pass
To identify potential TorchDynamo compilation entry points, this pass traverses the AST and analyzes each node. A node is tagged when a function or an nn.Module object wrapped or decorated with torch.compile is encountered. If no occurrences of torch.compile are found in the codebase, GraphMend compilation is skipped. Otherwise, the tagged nodes serve as entry points for subsequent passes.
Graph Break Type Analysis Pass.
We implement a compound analysis pass on UniiR to identify potential graph-break locations, following the procedure in Algorithm 1. Starting from each Dynamo entry point, the pass traverses the CFG to locate program points that may induce graph breaks. For conditional branches, the pass inspects the subexpressions of the condition. If any involve Torch attributes known to be dynamic (e.g., .sum()), the branch is marked as dynamic control flow. Otherwise, a use-def analysis over the symbol table and CFG determines whether the condition depends on model inputs or tensor values; if so, it is also tagged as dynamic. In addition, side-effecting operations such as print and logger.* calls are detected and annotated as logging-induced graph breaks as in Algorithm 1. These annotations guide the subsequent transformation passes that eliminate graph breaks.
AST Transformation Pass
After identifying graph break locations, we modify the AST according to the graph break tags detected in the previous pass, after verifying the transformation is safe. Based on these tags, one of the following transformations is applied.
@@
Before performing code transformation to address data-dependent control flow that causes graph breaks, we first analyze each branch body to identify which variables are modified and which functions are invoked in the if_true and if_false branches. If both branches produce the same output variables, their results can first be computed separately using two different temporary variables, and then combined with a torch.where expression that selects values based on the condition. If both branches call the same function, the function is moved outside the branches and executed only once, after its input arguments are merged using torch.where.
The transformation proceeds by rewriting the AST to handle data-dependent control-flow graph breaks. The branch condition is first evaluated and stored as a predicate (pred). Each branch is then expressed as a dataflow computation producing temporary intermediate results, which are merged using torch.where(pred, e1, e2). After this reconciliation step, we handle common function calls ensuring that both their inputs and side-effect-free semantics are preserved. Finally, the original if and else nodes are removed, and their child nodes are reconnected to the parent, yielding a transformed AST in which both execution paths are represented explicitly as dataflow expressions. This process, illustrated in Figure 7, eliminates data-dependent control-flow constructs and enables predicated, differentiable execution compatible with graph compilation.
Table 1: Models in benchmark suite
biogpt Luo et al. (2022) 2 logger calls 408
blenderbot-400M-distill Meta AI (2020) 3 logger calls 489
flan-t5-large Chung et al. (2022) 3 logger calls 2130
longformer-base-409 Beltagy et al. (2020) 5 logger calls, tensor.item() call 913
moe-minicpm-x4-base babybirdprd (2025) 15 Dynamic shape operator 280
Phi-4-mini-instruct Microsoft (2025) 5 Dynamic control flow 1541
Qwen-Audio-Chat Chu et al. (2023) 2 Dynamic control flow 1647
tiny-random-PegasusForCausalLM Face (2025) 2 logger calls 64
Deferred Side Effect Transformation
Before applying this transformation, we perform analysis to ensure that print and logger statements correspond to the built-in implementations rather than user-overloaded versions. This prevents interference with user-defined behavior or unintended semantic changes. Once verified, side-effecting calls such as print or logger are rewritten into deferred buffer assignments. Specifically, the value to be printed/ logged is stored in a dedicated runtime buffer together with a predicate flag indicating whether the branch containing the statement was executed. After the FX graph finishes execution, the buffer is flushed: each stored message is printed or logged only if its flag is true. This keeps side-effecting operations in the right order and with the same behavior, but safely outside the traced region.
@@
5 Evaluation
Table 2: Graph break counts in the original model and fix rates achieved by applying GraphMend across the benchmark suite.
@@ 5 Evaluation
Experimental Setup
To implement our compiler technique, we build on Jaseci framework Mars et al. (2023); Mars (2025). Experiments are conducted on NVIDIA RTX 3090 and A40 GPUs using TorchInductor as the backend compiler.
Benchmark Suite
To construct our benchmark suite, we (1) randomly sample 65 Hugging Face Hugging Face (2025) models spanning diverse architectures and task categories; (2) test each model to identify whether it triggers FX graph breaks; and (3) retain all 8 models that exhibit graph breaks for evaluation. Table 1 summarizes the resulting suite, including the number and causes of graph breaks and the total operator count. A similar methodology was used in prior work Ansel et al. (2024).
Profiling Methodology
We profile the forward pass using PyTorch Profiler pyt (2021). We evaluate the original and graph-break–fixed versions of each model, running seven iterations with one cold start and six warm runs. The profiler records CPU, CUDA, and operator-level events, which are then analyzed using Chrome’s trace viewer.
5.1 GraphMend’s Ability to Fix Graph Breaks
First, we evaluate GraphMend’s effectiveness in fixing graph breaks across the benchmark suite. As shown in Table 2, GraphMend fully eliminates graph breaks in 6 models and resolves some graph breaks in 2 others. Graph breaks introduced by tensor.item() calls and dynamic-shape operators remain unfixed, as they cannot be resolved through our code transformations. For the performance analysis that follows, we report results on the 7 models where graph breaks were fully or partially eliminated, excluding moe-minicpm-x4-base.
5.2 Latency Improvement per Forward Pass Using GraphMend to Fix Graph Breaks
We measure the relative latency improvement as the percentage reduction in per-forward-pass latency after applying GraphMend fixes compared to the version with graph breaks. Results are reported for the models using PyTorch Profiler data from both cold and warm runs.
Cold Start Latency Improvement
Figure 8(a) shows results on RTX 3090 and A40 GPUs. Across models, we observe a 30–75% reduction in latency during cold runs. At cold start, each graph break introduces significant overhead as separate CUDA graphs must be recorded and cached. Eliminating these breaks reduces capture operations to a single (or fewer) CUDA graph, minimizing overhead and yielding larger per forward pass improvements. A detailed analysis appears in  § 5.4.
Steady-State Latency Improvement
Figure 8(b) shows steady state results on RTX 3090 and A40 GPUs, with 2.5–25% latency reductions across models. After the cold run, cached CUDA graphs eliminate capture overhead, but remaining graph breaks still cause GPU to CPU switches, adding latency. Removing these breaks with GraphMend eliminates such switches, improving forward pass performance.
Notably, tiny-random-PegasusForCausalLM shows over 25% improvement. Since this model is very small (64 operators), the fixed costs from graph breaks (e.g. host–device synchronization, and extra graph launches) take up a large portion of the runtime. Removing these breaks therefore gives a huge speedup per forward pass.
5.3 Relative Throughput Improvement by Fixing Graph Breaks
We measure the relative throughput improvement as the percentage increase in output tokens per second after applying GraphMend fixes compared to the version with graph breaks.
We measure relative throughput for all models on RTX 3090 and A40 GPUs. Figure 8(b) shows 5–8% improvements, consistent with the latency reductions in Section 5.2. As each forward pass becomes faster, overall tokens per second increase. The variation of improvement across models reflects how much a graph break affects the forward path: when the impact of a break is small, the corresponding throughput gain is also limited.
For example, in both Phi-4-mini-instruct and Qwen-Audio-Chat, the root cause of the breaks is dynamic control flow. We observe a 7.5%–8% improvement for Qwen-Audio-Chat, compared to 5%–6% for Phi-4-mini-instruct. The difference arises because the break in Phi-4-mini-instruct occurs early in the forward function, while in Qwen-Audio-Chat it occurs mid-function, where its impact is greater. Hence, a break’s location determines its effect, fixing some yields larger throughput gains than others.
5.4 Overhead Analysis of Graph Breaks
We analyze how fixing graph breaks with GraphMend reduces overhead using the Qwen Audio Chat model on an A40 GPU, which contains two graph breaks.
For a consistent comparison, all activity plots (Figures 11(a)11(d)) show the time interval between the first and second CUDA graph executions in the original model; the same interval is used for the fixed model.
@@ 5.4 Overhead Analysis of Graph Breaks
Figure 10(a) shows the profiler trace for the original model. Execution is split into 3 CUDA graphs due to the 2 graph breaks, with idle GPU periods between the first and second FX graph executions. This is clearly visible in Figure 11(a), GPU becomes idle while CPU handles control flow in eager mode, including a D2H memcpy triggered by dynamic control flow. This CPU fallback introduces an overhead.
After applying GraphMend replacing torch.where, the resulting trace (Figure 10(b)) shows one continuous CUDA graph. As confirmed in Figure 11(c), GPU activity remains uninterrupted and CPU fallbacks disappear. Thus, GraphMend eliminates the overhead of CPU–GPU context switching during steady-state runs.
@@
Cold runs introduce even larger overheads. In the original model (Figure 11(b)), each graph break triggers a separate CUDA Graph recording and caching, causing long GPU idle phases and heavy CPU activity. GPU idle time dominates as the CPU handles repeated captures. In contrast, the fixed model (Figure 11(d)) performs a single capture, minimizing GPU idle time and CPU work. Consequently, GraphMend yields larger overhead reductions in cold runs, explaining the substantial cold-start latency gains per forward pass shown in Figure 8(a).
5.5 GPU Runtime Kernel Statistics Analysis
We examine how fixing graph breaks affects GPU kernel execution and optimization. On Phi-4-mini-instruct model (A40 GPU), the profiler reports 404 kernels for the original model and 393 for the fixed one; a reduction reflecting increased fusion and fewer launches as larger regions are captured.
As shown in Figure 12, three small kernels in the original model merge into a single fused kernel after fixing. Kernel reordering also appears, as the unified CUDA Graph enables better packing and scheduling of independent regions, reducing stalls and improving locality.
Graph breaks previously forced separate captures, preventing cross-break fusion and limiting scheduling windows. Their removal allows broader fusion, redundant launch elimination, and more flexible reordering; reducing kernel gaps and boosting device utilization.
@@ 5.5 GPU Runtime Kernel Statistics Analysis
6 Related Work
Table 3: Comparison of GraphMend with related transformation frameworks.
Feature AutoGraph MagPy TensorSSA GraphMend
Moldovan et al. (2019) Zhang et al. (2024) Ma et al. (2024)
[…]
Fully compatible with PyTorch 2 pipeline $cellcolorlightred \cellcolorlightred \cellcolorlightgreen\checkmark$
Handling side effects compatible with PyTorch $cellcolorlightred \cellcolorlightred \cellcolorlightgreen\checkmark$
[…]
PyTorch has long sought to balance performance with Python’s flexibility. TorchScript’s tracing and scripting enabled ahead-of-time graph capture but struggled with dynamic control flow and unsupported constructs PyTorch-TorchScript Team (2025); PyTorch Team (2025). PyTorch 2 introduced TorchDynamo for bytecode-level graph capture Ansel et al. (2024); Paszke et al. (2019), yet data-dependent branches and Python side effects still cause frequent graph breaks. GraphMend mitigates these by transforming the Python AST before execution.
While AutoGraph Moldovan et al. (2019) generates computation graphs in TensorFlow through source-level transformations, its approach does not generalize to PyTorch 2, which captures graphs at the bytecode level. In contrast, GraphMend performs lightweight AST-level transformations integrated directly into the PyTorch 2 compilation pipeline, improving symbolic evaluation and graph completeness. MagPy Zhang et al. (2024) takes a different route by replacing TorchDynamo and reconstructing graphs from runtime traces, whereas GraphMend preserves the standard compilation flow with minimal overhead. TensorSSA Ma et al. (2024) introduces a custom SSA-based IR to manage tensor mutations and control flow but adds significant infrastructure and omits Python side effects handling. GraphMend instead applies predicate-based SSA-inspired rewrites directly on the Python AST, avoiding extra IR layers while staying fully compatible with PyTorch’s ecosystem.
More broadly, GraphMend’s source-level rewriting follows the tradition of semantics-preserving transformations in program optimization, as exemplified by Halide Ragan-Kelley et al. (2013), Lantern Wang et al. (2018), and DLVM Zhao et al. (2017). Like these systems, GraphMend leverages structural code analysis for performance gains but is uniquely tailored to PyTorch, enabling TorchDynamo to capture larger, uninterrupted graphs. General compiler infrastructures such as XLA Community (2023), TVM Chen et al. (2018), and MLIR Lattner et al. (2020) provide IR-based optimization, while Glow Rotem et al. (2018), TASO Jia et al. (2019), and ONNX Runtime Microsoft (2018) extend portability across hardware and frameworks. These systems assume large graphs are already available, whereas GraphMend restructures PyTorch code to maximize graph continuity so their optimizations can be applied more effectively.
7 Conclusion
We present GraphMend, a source-level compiler technique for PyTorch that systematically eliminates common FX graph breaks before execution. By restructuring code constructs such as dynamic control flow and Python side effects (e.g., I/O operations), GraphMend reduces fragmentation in the compilation pipeline and enables TorchDynamo to capture larger, unified graphs. Our evaluation across a diverse suite of Hugging Face models demonstrates that GraphMend removes nearly all fixable graph breaks and consistently improves runtime efficiency, achieving substantial reductions in latency and increases in throughput on modern GPUs.